OpenStudy (anonymous):

What is the equation of the line that is perpendicular to -x + y = 7 and passes through (-1, -1)?

OpenStudy (anonymous):

the slope of the line that is perpendicular to the line -x+y=7 is -1. -1=y+1/x+1 -x-y=2